# Non-toric brane webs, Calabi–Yau 3-folds, and 5d SCFTs

### v1

@@ 4.2. Geometric classification of consistent asymptotic webs of 5-branes with 7-branes
First assume that $W^{\mathrm{asym}}$ is of Type I, so that $(Y,D)$ is generic. Let $C$ be a general curve in $|L|$. Since $(Y,D)$ is generic, every reduced irreducible component of $C$ with negative self-intersection is necessarily an interior $(-1)$-curve. Hence, one can write $C=C^{\prime}+\sum_{i}a_{i}E_{i}$, where all irreducible components of $C^{\prime}$ have nonnegative self-intersection, $w_{i}\in\mathbb{Z}_{\geq 0}$, and $E_{i}$ are interior $(-1)$-curves. If $E_{i}\cap E_{j}\neq 0$, then $(E_{i}+E_{j})^{2}\geq 0$, and the curve $E_{i}+E_{j}$ can be deformed non-trivially in $Y$, in contradiction with the assumption that $C$ is general in $|L|$. Hence, the curves $E_{i}$ are disjoint. Since all irreducible components of $C^{\prime}$ have nonnegative self-intersection, $\mathcal{O}(C^{\prime})$ is nef by [42, Lemma 4.14]. On the other hand, we have $C\cdot E_{i}=C^{\prime}\cdot E_{i}-a_{i}>0$ for all $i$ since $W^{\mathrm{asym}}$ is consistent, and so $L=\mathcal{O}(C)$ is also nef. Moreover, we have
@@
If $C^{2}>0$, then $L=\mathcal{O}(C)$ is big and nef, and so, by [42, Theorem 4.12], we have the following possibilities:
@@
    $C=C^{\prime}+E$ for an interior $(-1)$-curve such that $C^{\prime}\cdot E=1$. In particular, we have $C\cdot E=0$, which contradicts the consistency of $W^{\mathrm{asym}}$.
@@
If $C^{2}=0$, then we necessarily have $a_{i}=0$ for all $i$, and so Lemma 4.6 i) also follows in this case. Finally, if $C^{2}=0$, Lemma 4.6 iii) follows by [42, Theorem 4.19]. The situation where $(Y,D)$ is of Type II can be treated similarly. ∎
@@
Using that the exceptional curves $E_{e,i}$ are $(-1)$-curves, it follows from the definition of $L$ in Equation (3.3) that we have
@@
In case i), using that $L^{2}>0$, $L$ is nef, and all connected components $C_{i}$ of $C$ satisfy $C_{i}^{2}\geq 0$ by Lemma 4.6, the Case III of the proof of [42, Theorem 4.12] implies that the general curve $C\in|L|$ is connected. Equation (4.7) follows from the adjunction formula for curves on surfaces and from the fact that $D$ is an anticanonical divisor on $Y$. In case ii), the result follows from [42, Theorem 4.19]. ∎
@@
If $L^{2}>0$, that is, in case i) of Lemma 4.8, we necessarily have $g=\frac{1}{2}(L^{2}-L\cdot D)+1\geq 0$ since the genus of a smooth connected curve is non-negative. The non-negative integer $g$ is the rank, that is, the dimension of the Coulomb branch of the corresponding 5d SCFT defined by the consistent asymptotic web $W^{\mathrm{asym}}$. This numerical constraint on webs has been discussed in [19, 33, 62, 94, 95] and is referred to as the “r-rule” in [94, 95]. However, contrary to general expectations in the physics literature, if $W^{\mathrm{asym}}$ is a supersymmetric but not necessarily consistent asymptotic web, then $g=\frac{1}{2}(L^{2}-L\cdot D)+1$ is not in general the rank of the 5d SCFT defined by $W^{\mathrm{asym}}$ and does not have to be nonnegative. Indeed, while $g=\frac{1}{2}(L^{2}-L\cdot D)+1$ is always the arithmetic genus of a curve $C\in|L|$, such a curve is not necessarily connected if the web is not consistent. In general, $C$ is a disjoint union of a smooth connected curve of genus $g^{\prime}\geq 0$ and of $k\geq 0$ smooth curves of genus 00, so that $g^{\prime}=g+k$, and the 5d SCFT defined by the supersymmetric web is obtained by adding $k$ free hypermultiplets to a rank $g^{\prime}$ 5d SCFT defined by a consistent web, as explained below Definition 4.5. We present a detailed example of this phenomenon in Example 5.9.
@@
    If $L^{2}>0$, then $W^{\mathrm{asym}}$ is consistent if and only if $L$ is nef, the linear system $|L|$ has no fixed component, and, for every curve $E$ in $Y$ whose strict transform in a corner blow-up of $Y$ is an interior $(-1)$-curve or an internal $(-2)$-curve, we have $L\cdot E>0$.
@@
    If $L^{2}=0$, then $W^{\mathrm{asym}}$ is consistent if and only if $L$ is nef, $|L|\neq 0$ and the linear system $|L|$ has no fixed component.
@@
To prove the “if” direction of i), we first note that, if $L^{2}>0$, $L$ is nef, and no component of $D$ is a fixed component of $|L|$, then the linear system $|L|$ is non-empty by [42, Lemma 4.10]. If $|L|$ has no fixed component, then, by [42, Theorem 4.12], $|L|$ has also no base point, except if $L\cdot\widetilde{D}=1$, in which case the unique smooth point $p\in D$ such that $L|_{D}=\mathcal{O}_{D}(p)$ is the unique base point of $|L|$. If $|L|$ is base point free, then the general element of $|\widetilde{L}|$ is smooth by Bertini theorem [59, Corollary III.10.9]. If $|L|$ has a base point $p$, then, since $L\cdot D=1$, $p$ is a smooth point of every element of $|L|$, and after blowing up $p$, we obtain a base point free linear system, whose general element is smooth by Bertini theorem [59, Corollary III.10.9]. In any case, the general curve $C\in|L|$ is smooth. Since $|L|$ has no fixed component, it follows that every connected component $C_{i}$ of $C$ satisfies $C_{i}^{2}\geq 0$, and so $W^{\mathrm{asym}}$ is consistent. Finally, the “if” direction in ii) follows from [42, Theorem 4.19]. ∎
Finally, we are ready to provide a geometric classification of consistent asymptotic webs $W^{\mathrm{asym}}$, up to Hanany–Witten moves. To do this, denoting by $(Y,D,L)$ the associated log Calabi–Yau surface with line bundle, we analyze the image of the contraction map $c$ on $Y$, which contracts all curves that have intersection number 00 with $L$. The existence of such a contraction map $c$ is ensured by the nefness of $L$ shown in Lemma 4.6 and the basepoint free theorem in birational geometry [69, Theorem 3.3]. By Lemma 4.6, we have one of the three following cases to analyze:
@@
$\mathbf{EllLogCY}$: the set of all log Calabi–Yau surfaces $(Y,D)$, with a line bundle $L$, such that $|L|$ defines an elliptic fibration $Y\to\mathbb{P}^{1}$, up to deformations preserving the elliptic fibration. These log Calabi–Yau surfaces can be explicitly classified. The fibers of $c$ away from $D$ have at worst $I_{1}$ singular fibers for general deformations. On the other hand, $D$ is a cycle of rational curves. Since $Y$ is a rational elliptic surface, if $Q$ denotes the number of $I_{1}$ fibers away from $D$, then $D$ is a cycle of $12-Q$ irreducible rational curves. One can show that we necessarily have $3\leq Q\leq 11$. Moreover, for every such $3\leq Q\leq 11$ with $Q\neq 4$, there exists a unique deformation class of rational elliptic surfaces, and if $Q=4$, there exist exactly two classes of deformation of rational surfaces [87]. If $L=\mathcal{O}(kF)$, where $k\geq 1$ and $F$ a general fiber of $c$, the corresponding 5d SCFT is the rank $k$ version of the $E_{Q-3}$ theory if $n\neq 4$, or of the $E_{1}$ and $\widetilde{E}_{1}$ theories if $Q=4$ [29, 32, 34, 83, 90, 97].
@@
Injectivity follows as in the proof of Theorem 3.12. Indeed, if two webs of 5-branes with 7-branes define two toric models of the same log Calabi–Yau surface with line bundle up to corner blow-ups and blow-downs, then, by [55, Proposition 3.27] these two toric models are connected by a sequence of corner blow-ups/blow-downs, and so the corresponding webs are connected by a sequence of Hanany–Witten moves.
To prove the surjectivity of $F_{I}$ or $F_{II,+}$, let $(Y,D,L)$ be a smooth resolution of $(Y^{\mathrm{pol}},D^{\mathrm{pol}},L^{\mathrm{pol}})\in\mathbf{PolLogCY}$ or $(Y^{\mathrm{pol}},L^{\mathrm{pol}})\in\mathbf{PolCY^{cusp}}$. Modulo additional corner blow-ups, we can assume that $D$ has at least three irreducible components. Since $L^{\mathrm{pol}}$ is ample, the line bundle $L$ is big and nef, and so it follows from the proof of [37, Theorem 5.4] that there exists a toric model $p:(Y,D)\rightarrow(\overline{Y},\overline{D})$, with interior exceptional curves $E_{i}$, and an ample line bundle $\overline{L}$ on $\overline{Y}$ such that $L=p^{\star}\overline{L}\otimes\mathcal{O}(-\sum_{i}a_{i}E_{i})$ with $a_{i}\in\mathbb{Z}_{\geq 0}$. Let $W^{\mathrm{asym}}$ be the web of 5-branes with 7-branes obtained from the web of 5-branes associated to the polarized toric surface $(\overline{Y},\overline{D},\overline{L})$ by ending $a_{i}$ 5-branes on a $(p_{i},q_{i})$ 7-brane, where $\mathbb{R}_{\geq 0}(p_{i},q_{i})$ is the ray of $\overline{W}$ corresponding to the toric divisor of $\overline{Y}$ containing the point $p(E_{i})$. Then, by construction, we have $F_{I}(W^{\mathrm{asym}})=(Y^{\mathrm{pol}},D^{\mathrm{pol}},L^{\mathrm{pol}})$ or $F_{II,+}(W^{\mathrm{asym}})=(Y^{\mathrm{pol}},L^{\mathrm{pol}})$. Finally, it remains to prove the surjectivity of $F_{II,0}$. For $(Y,D,L)\in\mathbf{EllLogCY}$, we have $L=\mathcal{O}(kD)$ for some $k\in\mathbb{Z}_{\geq 1}$. Hence, the analog of [37, Proposition 1.5] holds, and so one can run the argument in the proof of [37, Theorem 5.4], and conclude the surjectivity of $F_{II,0}$ as done above for $F_{I}$ and $F_{II,+}$. ∎
@@ 4.3. Consistent webs of 5-branes with 7-branes and birational geometry of degenerations
Let $E$ be an interior $(-1)$-curve in an irreducible component $Y_{0}^{v}$ of $\mathcal{Y}_{0}$, corresponding to a vertex $v$ of $W$. Assume that the intersection point $p=E\cap\partial Y_{0}^{v}$ is not contained in the $\mathcal{D}_{0}$. Thus, there exists another component $Y_{0}^{v^{\prime}}$ of $\mathcal{Y}_{0}$, corresponding to a vertex $v^{\prime}$ of $W$, such that $p\in Y_{0}^{v}\cap Y_{0}^{v^{\prime}}$. An M1 flop of $E$ applied to $\mathcal{Y}_{0}$ is the blow-down of $E$ in $Y_{0}^{v}$, followed by the blow-up of the point $p$ in $Y_{0}^{v^{\prime}}$. We denote by $E^{\prime}$ the exceptional curve in the blow-up of $Y_{0}^{v^{\prime}}$.
An M1 flop of $E$ applied to $\mathcal{Y}_{0}$ uniquely extends to a birational modification of $\mathcal{Y}$, which is the identity away from $\mathcal{Y}_{0}$. Indeed, the normal bundle of $E\simeq\mathbb{P}^{1}$ in $Y$ is $\mathcal{O}_{\mathbb{P}^{1}}(-1)\oplus\mathcal{O}_{\mathbb{P}^{1}}(-1)$ and the M1 flop on $\mathcal{Y}$ is the Atiyah flop corresponding to $E$ – see [15]. We refer to this Atiyah flop, as the M1 flop of $E$ applied to $\mathcal{Y}$, following the terminology of [6]. Since an M1 flop is an isomorphism on codimension $2$, denoting by $\mathcal{Y}^{\prime}$ the resulting 3-fold obtained from $\mathcal{Y}$ by an M1 flop, we have a natural isomorphism of the Picard groups $\mathrm{Pic}(\mathcal{Y})\simeq\mathrm{Pic}(\mathcal{Y}^{\prime})$. Hence, the line bundle $\mathcal{L}$ corresponds to a line bundle $\mathcal{L}^{\prime}$ on $\mathcal{Y}^{\prime}$. A key feature of an M1 flop is that it changes the intersection numbers by:
@@
Since $W$ is consistent, there exists a family of curves $(C_{t})_{t}$ in $\mathcal{Y}$, such that the central fiber $C_{0}\in|\mathcal{L}_{0}|$ in $\mathcal{Y}_{0}$ does not contain any irreducible component of $\mathcal{D}_{0}$ or of the double locus of $\mathcal{Y}_{0}$. Assume that $\mathcal{L}$ is not nef, and let $E$ be an irreducible curve such that $\mathcal{L}\cdot E<0$. Since $C_{0}$ does not contain any irreducible component of $\mathcal{D}_{0}$ or of the double locus of $\mathcal{Y}_{0}$, $E$ must be an interior $(-1)$-curve by [92, Lemma 1.4]. Applying an M1 flop of $E$ and denoting by $\mathcal{Y}^{\prime}$ the resulting 3-fold with line bundle $\mathcal{L}^{\prime}$ as before, we get a (-1)-curve $E^{\prime}$ in $\mathcal{Y}^{\prime}$, which by (4.12) has positive intersection with $\mathcal{L}^{\prime}$. The resulting curve $C_{0}^{\prime}$ after the flop still does not contain any irreducible component of the double locus of $\mathcal{Y}_{0}^{\prime}$ or of $\mathcal{D}_{0}^{\prime}$. On the other hand, denoting by $D_{E}$ the irreducible component of the double locus of $\mathcal{Y}_{0}$ intersecting $E$, it follows from [92, Lemma 1.1] that

### v2

@@ 4.2. Geometric classification of consistent asymptotic webs of 5-branes with 7-branes
FFirst assume that $W^{\mathrm{asym}}$ is of Type I, so that $(Y,D)$ is generic. Let $C$ be a general curve in $|L|$. Since $(Y,D)$ is generic, every reduced irreducible component of $C$ with negative self-intersection is necessarily an interior (1)(-1)-curve. Hence, one can write $C=C^{\prime}+\sum_{i}a_{i}E_{i}$, where all irreducible components of $C^{\prime}$ have nonnegative self-intersection, $w_{i}\in\mathbb{Z}_{\geq 0}$, and $E_{i}$ are interior (1)(-1)-curves. If $E_{i}\cap E_{j}\neq 0$, then $(E_{i}+E_{j})^{2}\geq 0$, and the curve $E_{i}+E_{j}$ can be deformed non-trivially in $Y$, in contradiction with the assumption that $C$ is general in $|L|$. Hence, the curves $E_{i}$ are disjoint. Since all irreducible components of $C^{\prime}$ have nonnegative self-intersection, $\mathcal{O}(C^{\prime})$ is nef by [45, Lemma 4.14]. On the other hand, we have $C\cdot E_{i}=C^{\prime}\cdot E_{i}-a_{i}>0$ for all $i$ since $W^{\mathrm{asym}}$ is consistent, and so $L=\mathcal{O}(C)$ is also nef. Moreover, we have
@@
If $C^{2}>0$, then $L=\mathcal{O}(C)$ is big and nef, and so, by [45, Theorem 4.12], we have the following possibilities:
@@
    $C=C^{\prime}+E$ for an interior (1)(-1)-curve such that $C^{\prime}\cdot E=1$. In particular, we have $C\cdot E=0$, which contradicts the consistency of $W^{\mathrm{asym}}$.
@@
If $C^{2}=0$, then we necessarily have $a_{i}=0$ for all $i$, and so Lemma 4.6 i) also follows in this case. Finally, if $C^{2}=0$, Lemma 4.6 iii) follows by [45, Theorem 4.19]. Finally, the situation where $(Y,D)$ is of Type II can be treated similarly. ∎
@@
Using that the exceptional curves $E_{e,i}$ are (1)(-1)-curves, it follows from the definition of $L$ in Equation (3.3) that we have
@@
In case i), using that $^{2}>0, LL$ is nef, and all connected components $C_{i}$ of $C$ satisfy $C_{i}^{2}\geq 0$ by Lemma 4.6, the Case III of the proof of [45, Theorem 4.12] implies that the general curve $C\in|L|$ is connected. Equation (4.7) follows from the adjunction formula for curves on surfaces and from the fact that $D$ is an anticanonical divisor on $Y$. In case ii), the result follows from [45, Theorem 4.19]. ∎
@@
If $L^{2}>0$, that is, in case i) of Lemma 4.8, we necessarily have $g=\frac{1}{2}(L^{2}-L\cdot D)+1\geq 0$ since the genus of a smooth connected curve is non-negative. The non-negative integer $g$ is the rank, that is, the dimension of the Coulomb branch of the corresponding 5d SCFT defined by the consistent asymptotic web $W^{\mathrm{asym}}$. This numerical constraint on webs has been discussed in [22, 36, 65, 97, 98] and is referred to as the “r-rule” in [97, 98]. However, contrary to general expectations in the physics literature, if $W^{\mathrm{asym}}$ is a supersymmetric but not necessarily consistent asymptotic web, then $g=\frac{1}{2}(L^{2}-L\cdot D)+1$ is not in general the rank of the 5d SCFT defined by $W^{\mathrm{asym}}$ and does not have to be nonnegative. Indeed, while $g=\frac{1}{2}(L^{2}-L\cdot D)+1$ is always the arithmetic genus of a curve $C\in|L|$, such a curve is not necessarily connected if the web is not consistent. In general, $C$ is a disjoint union of a smooth connected curve of genus $g^{\prime}\geq 0$ and of $k\geq 0$ smooth curves of genus 0, so that $g^{\prime}=g+k$, and the 5d SCFT defined by the supersymmetric web is obtained by adding $k$ free hypermultiplets to a rank $g^{\prime}$ 5d SCFT defined by a consistent web, as explained below Definition 4.5. We present a detailed example of this phenomenon in Example 5.10.
@@
    If $L^{2}>0$, then $W^{\mathrm{asym}}$ is consistent if and only if $L$ is nef, the linear system $|L|$ has no fixed component, and, for every curve $E$ in $Y$ whose strict transform in a corner blow-up of $Y$ is an interior (1)(-1)-curve or an internal (2)(-2)-curve, we have $L\cdot E>0$.
@@
    If $L^{2}=0$, then $W^{\mathrm{asym}}$ is consistent if and only if $L$ is nef and the linear system $|L|$ is non-empty and has no fixed component.
@@
[…]
Finally, we are ready to provide a geometric classification of consistent asymptotic webs $W^{\mathrm{asym}}$, up to Hanany–Witten moves. To do this, denoting by $(Y,D,L)$ the associated log Calabi–Yau surface with line bundle, we analyze the image of the contraction map $c$ on $Y$, which contracts all curves that have intersection number 0 with $L$. The existence of such a contraction map $c$ is ensured by the nefness of $L$ shown in Lemma 4.6 and the basepoint free theorem in birational geometry [72, Theorem 3.3]. By Lemma 4.6, we have one of the three following cases to analyze:
@@
$\mathbf{EllLogCY}$: the set of all log Calabi–Yau surfaces $(Y,D)$, with a line bundle $L$, such that $|L|$ defines an elliptic fibration $Y\to\mathbb{P}^{1}$, up to deformations preserving the elliptic fibration. These log Calabi–Yau surfaces can be explicitly classified. The fibers of $c$ away from $D$ have at worst $I_{1}$ singular fibers for general deformations. On the other hand, $D$ is a cycle of rational curves. Since $Y$ is a rational elliptic surface, if $Q$ denotes the number of $I_{1}$ fibers away from $D$, then $D$ is a cycle of 12Q12-Q irreducible rational curves. One can show that we necessarily have $3\leq Q\leq 11$. Moreover, for every such $3\leq Q\leq 11$ with $Q\neq 4$, there exists a unique deformation class of rational elliptic surfaces, and if $Q=4$, there exist exactly two classes of deformation of rational surfaces [90]. If $L=\mathcal{O}(kF)$, where $k\geq 1$ and $F$ a general fiber of $c$, the corresponding 5d SCFT is the rank $k$ version of the $E_{Q-3}$ theory if $n\neq 4$, or of the $E_{1}$ and $\widetilde{E}_{1}$ theories if $Q=4$ [32, 35, 37, 86, 93, 100].
@@
Injectivity follows as in the proof of Theorem 3.12. Indeed, if two webs of 5-branes with 7-branes define two toric models of the same log Calabi–Yau surface with line bundle up to corner blow-ups and blow-downs, then, by [58, Proposition 3.27] these two toric models are connected by a sequence of corner blow-ups/blow-downs, and so the corresponding webs are connected by a sequence of Hanany–Witten moves.
To prove the surjectivity of $F_{I}$ or $F_{II,+}$, let $(Y,D,L)$ be a smooth resolution of $(Y^{\mathrm{pol}},D^{\mathrm{pol}},L^{\mathrm{pol}})\in\mathbf{PolLogCY}$ or $(Y^{\mathrm{pol}},L^{\mathrm{pol}})\in\mathbf{PolCY^{cusp}}$. Modulo additional corner blow-ups, we can assume that $D$ has at least three irreducible components. Since $L^{\mathrm{pol}}$ is ample, the line bundle $L$ is big and nef, and so it follows from the proof of [40, Theorem 5.4] that there exists a toric model $p:(Y,D)\rightarrow(\overline{Y},\overline{D})$, with interior exceptional curves $E_{i}$, and an ample line bundle $\overline{L}$ on $\overline{Y}$ such that $L=p^{\star}\overline{L}\otimes\mathcal{O}(-\sum_{i}a_{i}E_{i})$ with $a_{i}\in\mathbb{Z}_{\geq 0}$. Let $W^{\mathrm{asym}}$ be the web of 5-branes with 7-branes obtained from the web of 5-branes associated to the polarized toric surface $(\overline{Y},\overline{D},\overline{L})$ by ending $a_{i}$ 5-branes on a $(p_{i},q_{i})$ 7-brane, where $\mathbb{R}_{\geq 0}(p_{i},q_{i})$ is the ray of $\overline{W}$ corresponding to the toric divisor of $\overline{Y}$ containing the point $p(E_{i})$. Then, by construction, we have $F_{I}(W^{\mathrm{asym}})=(Y^{\mathrm{pol}},D^{\mathrm{pol}},L^{\mathrm{pol}})$ or $F_{II,+}(W^{\mathrm{asym}})=(Y^{\mathrm{pol}},L^{\mathrm{pol}})$. Finally, it remains to prove the surjectivity of $F_{II,0}$. For $(Y,D,L)\in\mathbf{EllLogCY}$, we have $L=\mathcal{O}(kD)$ for some $k\in\mathbb{Z}_{\geq 1}$. Hence, the analog of [40, Proposition 1.5] holds, and so one can run the argument in the proof of [40, Theorem 5.4], and conclude the surjectivity of $F_{II,0}$ as done above for $F_{I}$ and $F_{II,+}$. ∎
@@ 4.3. Consistent webs of 5-branes with 7-branes and birational geometry of degenerations
Let $E$ be an interior (1)(-1)-curve in an irreducible component $Y_{0}^{v}$ of $\mathcal{Y}_{0}$, corresponding to a vertex $v$ of $W$. Assume that the intersection point $p=E\cap\partial Y_{0}^{v}$ is not contained in the $\mathcal{D}_{0}$. Thus, there exists another component $Y_{0}^{v^{\prime}}$ of $\mathcal{Y}_{0}$, corresponding to a vertex $v^{\prime}$ of $W$, such that $p\in Y_{0}^{v}\cap Y_{0}^{v^{\prime}}$. An M1 flop of $E$ applied to $\mathcal{Y}_{0}$ is the blow-down of $E$ in $Y_{0}^{v}$, followed by the blow-up of the point $p$ in $Y_{0}^{v^{\prime}}$. We denote by $E^{\prime}$ the exceptional curve in the blow-up of $Y_{0}^{v^{\prime}}$.
An M1 flop of $E$ applied to $\mathcal{Y}_{0}$ uniquely extends to a birational modification of $\mathcal{Y}$, which is the identity away from $\mathcal{Y}_{0}$. Indeed, the normal bundle of $E\simeq\mathbb{P}^{1}$ in $\mathcal{Y}$ is $\mathcal{O}_{\mathbb{P}^{1}}(-1)\oplus\mathcal{O}_{\mathbb{P}^{1}}(-1)$ and the M1 flop on $\mathcal{Y}$ is the Atiyah flop corresponding to $E$ – see [18]. We refer to this Atiyah flop as the M1 flop of $E$ applied to $\mathcal{Y}$, following the terminology of [6]. Since an M1 flop is an isomorphism on codimension $2$, denoting by $\mathcal{Y}^{\prime}$ the resulting 3-fold obtained from $\mathcal{Y}$ by an M1 flop, we have a natural isomorphism of the Picard groups $\mathrm{Pic}(\mathcal{Y})\simeq\mathrm{Pic}(\mathcal{Y}^{\prime})$. Hence, the line bundle $\mathcal{L}$ corresponds to a line bundle $\mathcal{L}^{\prime}$ on $\mathcal{Y}^{\prime}$. A key feature of an M1 flop is that it changes the intersection numbers by:
@@
Since $W$ is consistent, there exists a family of curves $(C_{t})_{t}$ in $\mathcal{Y}$, such that the central fiber $C_{0}\in|\mathcal{L}_{0}|$ in $\mathcal{Y}_{0}$ does not contain any irreducible component of $\mathcal{D}_{0}$ or of the double locus of $\mathcal{Y}_{0}$. Assume that $\mathcal{L}$ is not nef, and let $E$ be an irreducible curve such that $\mathcal{L}\cdot E<0$. Since $C_{0}$ does not contain any irreducible component of $\mathcal{D}_{0}$ or of the double locus of $\mathcal{Y}_{0}$, $E$ must be an interior (1)(-1)-curve by [95, Lemma 1.4]. Applying an M1 flop of $E$ and denoting by $\mathcal{Y}^{\prime}$ the resulting 3-fold with line bundle $\mathcal{L}^{\prime}$ as before, we get a (-1)-curve $E^{\prime}$ in $\mathcal{Y}^{\prime}$, which by (4.12) has positive intersection with $\mathcal{L}^{\prime}$. The resulting curve $C_{0}^{\prime}$ after the flop still does not contain any irreducible component of the double locus of $\mathcal{Y}_{0}^{\prime}$ or of $\mathcal{D}_{0}^{\prime}$. On the other hand, denoting by $D_{E}$ the irreducible component of the double locus of $\mathcal{Y}_{0}$ intersecting $E$, it follows from [95, Lemma 1.1] that